# An explicit form of Ingham’s zero density estimate

### v1

@@ 2.1. The zero density estimate
Let $T\geq H\geq H_{0}\geq 1002$, $T\geq T_{0}\geq\max\left\{e^{2d},e^{e}\right\}$ for $d>0$, $0<\delta\leq\frac{1}{2}\log{(hT_{0})}$ for $h\geq 10^{9}/T_{0}$, and $\kappa>0$. Then
$N(\sigma,T)-N(\sigma,H)$ $\leq T^{\frac{3(1-\sigma)}{2-\sigma}}\left(\frac{\mathfrak{A}_{1}}{2\pi d}\left(\log{(hT)}\right)^{\frac{2\sigma}{2-\sigma}}\left(\log{T}\right)^{\frac{6-5\sigma}{2-\sigma}}+\frac{\mathfrak{A}_{2}}{2\pi d}\left(\log{(hT)}\right)^{\frac{2(2\sigma-1)}{2-\sigma}}\log{T}\right)$
@@
×$\times\max\left\{1,K_{3}^{-\frac{3d}{\log{T_{0}}}}\right\}\max\left\{1,K_{1}^{\frac{3d}{\log{T_{0}}}}\right\}\max\left\{L_{0}^{\frac{2}{2-\sigma_{1}}},L_{0}^{\frac{2}{2-\sigma_{2}}}\right\}\times$
×$\times\max\left\{K_{3}^{\frac{2\sigma_{1}-1}{2-\sigma_{1}}},K_{3}^{\frac{2\sigma_{2}-1}{2-\sigma_{2}}}\right\}\max\left\{K_{1}^{\frac{3(1-\sigma_{1})}{2-\sigma_{1}}},K_{1}^{\frac{3(1-\sigma_{2})}{2-\sigma_{2}}}\right\}$
@@
×$\times\max\left\{1,K_{3}^{-\frac{3d}{\log{T_{0}}}}\right\}\max\left\{L_{0}^{\frac{2}{2-\sigma_{1}}},L_{0}^{\frac{2}{2-\sigma_{2}}}\right\}\max\left\{K_{3}^{\frac{2\sigma_{1}-1}{2-\sigma_{1}}},K_{3}^{\frac{2\sigma_{2}-1}{2-\sigma_{2}}}\right\}\max\left\{K_{2}^{\frac{3(1-\sigma_{1})}{2-\sigma_{1}}},K_{2}^{\frac{3(1-\sigma_{2})}{2-\sigma_{2}}}\right\},$
@@
$N(\sigma,T)$ $\leq\frac{1}{2\pi d}\left(\mathfrak{A}_{1}+\mathfrak{A}_{2}\cdot\left(\log{H_{\mathrm{RH}}}\right)^{-\frac{6(1-\sigma_{2})}{2-\sigma_{2}}}\right)T^{\frac{3(1-\sigma)}{2-\sigma}}\log^{3}{T}$
@@
$N(\sigma,T)\leq 5.36\cdot T^{\frac{3(1-\sigma)}{2-\sigma}}\log^{3}{T}+9.461\log^{2}{T}+167.8\log{T}$
@@
$N(T)\leq\frac{T}{2\pi}\log{\frac{T}{2\pi e}}+0.43\log{T}\leq 5.36\exp{\left(\frac{-0.56}{\frac{3}{2}-\frac{0.28}{\log{H_{\mathrm{RH}}}}}\right)}T\log^{3}{T}+9.461\log^{2}{T}+167.8\log{T}$ (11)
@@ 2.2. The fourth power moment estimate
Let $A_{0}\geq 10^{5}$. Then we have $\mathcal{M}_{2}(A_{0})\leq a_{1}(A_{0})$ and $\mathcal{M}_{2}(A_{0},T)\leq a_{2}(A_{0})T\log^{4}{T}$ for every $T\geq A_{0}$, where
$a_{1}(A_{0})\mathrel{\mathop{:}}=\frac{1}{2\pi^{2}}A_{0}\log^{4}\frac{A_{0}}{2}+48.801A_{0}\log^{3}\frac{A_{0}}{2}+3.0592\cdot 10^{10}$
@@
$a_{2}(A_{0})\mathrel{\mathop{:}}=\frac{1}{2\pi^{2}}\left(1-\frac{\log{2}}{\log{A_{0}}}\right)^{4}+\frac{48.801}{\log{A_{0}}}\left(1-\frac{\log{2}}{\log{A_{0}}}\right)^{3}+\frac{2.1817\cdot 10^{10}}{A_{0}\log^{4}{A_{0}}}.$
@@
The first estimate follows by the second inequality in (9) of Corollary 2. Because $\mathcal{M}_{2}\left(A_{0},T\right)\leq\mathcal{M}_{2}\left(10^{5},T\right)$, we have $\mathcal{M}_{2}(A_{0},T)\leq a_{2}(T)T\log^{4}{T}$ by (18), see the proof of Corollary 2. Derivative analysis shows that $a_{2}(T)$ is a decreasing function for $T\geq 10^{5}$, implying that $a_{2}(T)\leq a_{2}(A_{0})$. This gives the second estimate. ∎
@@
Assume $T\geq T_{0}^{\prime\prime}=T_{0}\geq T_{0}^{\prime}=3\cdot 10^{3}$, and also that the conditions from Theorem 2 are satisfied. Note that $\mathcal{M}_{2}\left(T_{0}^{\prime\prime},2T_{0}^{\prime\prime}\right)\leq\mathcal{M}_{2}\left(T_{0}^{\prime},2T_{0}^{\prime\prime}\right)$. By (14) and (17) we have
$\mathcal{M}_{2}\left(T_{0}^{\prime\prime},T\right)\leq\frac{1}{2\pi^{2}}T\log^{4}{\frac{T}{2}}+\mathfrak{F}_{1}\left(T_{0}^{\prime\prime},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)T\log^{3}{\frac{T}{2}}+\mathcal{C}_{2}\left(T_{0}^{\prime\prime}\right),$
where
@@
We are taking $T_{0}^{\prime\prime}=10^{5}$. Finding a global minimum for $\mathfrak{F}_{1}\left(T_{0}^{\prime\prime},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)$ is a computationally very complex problem, since we need to optimize 11 parameters that satisfy the conditions (2)–(4) from Theorem 2. To simplify the process, we used randomization; we generate $50$ trials where for each trial we randomly choose $100$ sets of values for these parameters, then we select an element from this set such that $\mathfrak{F}_{1}\left(T_{0}^{\prime\prime},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)$ is minimal, and finally the “minimal” set from all trials. The values we obtained with such a method are $\bm{c}=(-0.938,0.194)$, $\bm{\sigma}=(0.368,(0.718,0.435),(0.629,0.3654))$, $\bm{\mathfrak{a}}=(1.99,1.44)$, and $\bm{\mathfrak{b}}=(1.33,1.45)$, giving $\mathfrak{F}_{1}\left(T_{0}^{\prime\prime},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)\leq 48.801$. The conditions (5) and (6) from Theorem 2 are clearly satisfied by this choice of parameters. Trivially, $\mathcal{C}_{2}\left(T_{0}^{\prime\prime}\right)\leq 2.1817\cdot 10^{10}$. Therefore,
@@
$\mathcal{M}_{2}\left(10^{5},T\right)\leq\frac{1}{2\pi^{2}}T\log^{4}{\frac{T}{2}}+48.801\cdot T\log^{3}{\frac{T}{2}}+2.1817\cdot 10^{10}$ (18)
for $T\geq 10^{5}$. The second inequality in (9) now easily follows from (18) since $\mathcal{M}_{2}\left(10^{5}\right)\leq 8.775\cdot 10^{9}$ by (8). The first inequality in (9) follows from (15) since $\mathcal{M}_{2}(T)\geq\mathcal{M}_{2}\left(10^{5},T\right)$. The proof of Corollary 2 is thus complete. ∎
@@
$\mathcal{M}_{1}(T)\leq\left(1+\frac{0.548}{\log{T_{0}}}\right)T\log{T},\quad\mathcal{M}_{2}(T)\leq\left(\frac{1}{2\pi^{2}}+\frac{241}{\log{T_{0}}}\right)T\log^{4}{T}$
for $T\geq T_{0}\geq 10^{5}$ by [DZA22, Theorem 3.4] for the first inequality, and by Corollary 2 and derivative analysis for the second inequality. The result now follows by the CBS inequality since then $\mathcal{M}_{2/3}(T)\leq\left(\mathcal{M}_{1}(T)\mathcal{M}_{2}(T)\right)^{1/2}$. ∎
@@ 3. Gabriel’s convexity theorem and applications
Let 1/$21/2\leq\alpha<1<\beta\leq 3/$ 2, $T\geq H>1$, $\kappa>0$, $A\geq 0$, and also let $\lambda$ and $\mu$ be arbitrary positive numbers. Then
@@
Let $T\geq T_{0}\geq 2$ and $11/2\leq\alpha<\beta<$ 1. Assume that $f(z)$ is a holomorphic function in the strip $x=\Re\{z\}\in(\alpha,\beta)$ with continuous extension to the boundary, such that the conditions (2) and (3) from Theorem 4 are satisfied. Assume also that
@@ 4. Proof of Theorem 1
Let $T>0$, $X\geq X_{0}\geq 10^{9}$, and $\delta>0$. Then
$F_{X}\left(1+\frac{\delta}{\log X},0,T,\frac{1}{2}\right)\leq\left(C_{3}(X_{0},\delta)+\frac{C_{4}(X_{0},\delta)\left(T+\pi m_{0}\right)}{X}\right)\log^{2}{X},$
@@
$C_{3}(X_{0},\delta)\mathrel{\mathop{:}}=\frac{\pi m_{0}b_{4}}{2\delta}\left(1+\frac{2\delta}{\log X_{0}}\right)^{2}\exp\left(\frac{2\delta\gamma}{\log X_{0}}\right)$
$C_{4}(X_{0},\delta)\mathrel{\mathop{:}}=\frac{b_{4}}{5\delta e^{\delta}}\left(1+\frac{\delta}{\log X_{0}}\right)^{2}+\frac{b_{3}e^{-2\delta}}{\log^{2}X_{0}},$
with $b_{3}\mathrel{\mathop{:}}=0.605$, $b_{4}\mathrel{\mathop{:}}=0.529$, $m_{0}$ as in Lemma 5, and $\gamma$ is the Euler–Mascheroni constant.
@@
See [KLN18, Lemma 4.3]. ∎
@@
Let $T\geq H\geq 2\pi m_{0}$, $T\geq T_{0}>0$ and $X=hT$ for $h\geq 10^{9}/T_{0}$. Then
$-\int_{H}^{T}\log{\left|h_{X}\left(\frac{3}{2}+\mathrm{i}t\right)\right|}\mathrm{d}t\leq C_{5}(hT_{0},h),$
@@
with $m_{0}$ as in Lemma 5 and $\gamma$ is the Euler–Mascheroni constant.
@@
$-\log{\left|h_{X}\left(\frac{3}{2}+\mathrm{i}t\right)\right|}\leq-\frac{\log{\left(1-b_{6}(hT_{0},1/2)^{2}\right)}}{b_{6}(hT_{0},1/2)^{2}}\left|f_{X}\left(\frac{3}{2}+\mathrm{i}t\right)\right|^{2}.$
@@
$-\frac{2\pi}{\log{2}}\log{\left(1-9\left(1+\frac{2+\gamma+\frac{7}{36u}}{\log{u}}\right)^{2}\frac{\log^{2}{u}}{u}\right)}$ (34)
and $\gamma$ is the Euler–Mascheroni constant.
@@
Let $T\geq H\geq H_{0}>1$, $T\geq T_{0}\geq e^{e}$, and $\kappa>0$. Also, let $X=hT$ for $h\geq 10^{9}/T_{0}$, and let $w_{1}=w_{1}\left(\sigma,1/2,1+\delta/\log{X}\right)$ and $w_{2}=w_{2}\left(\sigma,1/2,1+\delta/\log{X}\right)$ for $0<\delta\leq\frac{1}{2}\log{(hT_{0})}$ and with $w_{1}$ and $w_{2}$ as in (21). Then
$F_{X}\left(\sigma,H,T,\frac{3}{4}w_{1}+\frac{1}{2}w_{2}\right)$ $\leq L_{0}^{\frac{2}{2-\sigma}}K_{3}^{\frac{2\sigma-1}{2-\sigma}}\left(L_{1}\cdot\left(K_{1}T\right)^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{T}\right)^{\frac{4(1-\sigma)}{2-\sigma}}\left(\log{(hT)}\right)^{\frac{2\sigma}{2-\sigma}}\right.$
$\left.+L_{2}\cdot\left(K_{2}T\right)^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{(hT)}\right)^{\frac{2(2\sigma-1)}{2-\sigma}}\right)$
@@
$\int_{\frac{\kappa}{\mu}\left(\frac{A}{T}\right)^{2}}^{\infty}e^{-u}F_{X}\left(\beta,0,T\sqrt{\frac{\mu u}{\kappa}},\mu\right)\mathrm{d}u\leq K_{3}\log^{2}(hT)$

### v2

@@ 2.1. The zero density estimate
Let $\geq H\geq H_{0}\geq 1002, TT$$\geq T_{0}\geq\max\left\{e^{2d},e^{e}\right\}$ for $d>0$, $0<\delta\leq\frac{1}{20}\log{(hT_{0})}$ for $h\geq 3\cdot 10^{10}/T_{0}$, and $\kappa>0$. Then
$N(\sigma,T)-N(\sigma,H)$ $\leq T^{\frac{3(1-\sigma)}{2-\sigma}}\left(\frac{\mathfrak{A}_{1}}{2\pi d}\left(\log{(hT)}\right)^{\frac{1}{2-\sigma}}\left(\log{T}\right)^{\frac{6-5\sigma}{2-\sigma}}+\frac{\mathfrak{A}_{2}}{2\pi d}\left(\log{(hT)}\right)^{\frac{2\sigma-1}{2-\sigma}}\log{T}\right)$
@@
$\times\max\left\{1,K_{3}^{-\frac{3d}{\log{T_{0}}}}\right\}\max\left\{1,K_{1}^{\frac{3d}{\log{T_{0}}}}\right\}\max\left\{L_{0}^{\frac{2}{2-\sigma_{1}}},L_{0}^{\frac{2}{2-\sigma_{2}}}\right\}\times$
$\times\max\left\{K_{3}^{\frac{2\sigma_{1}-1}{2-\sigma_{1}}},K_{3}^{\frac{2\sigma_{2}-1}{2-\sigma_{2}}}\right\}\max\left\{K_{1}^{\frac{3(1-\sigma_{1})}{2-\sigma_{1}}},K_{1}^{\frac{3(1-\sigma_{2})}{2-\sigma_{2}}}\right\}$
@@
$\times\max\left\{1,K_{3}^{-\frac{3d}{\log{T_{0}}}}\right\}\max\left\{L_{0}^{\frac{2}{2-\sigma_{1}}},L_{0}^{\frac{2}{2-\sigma_{2}}}\right\}\max\left\{K_{3}^{\frac{2\sigma_{1}-1}{2-\sigma_{1}}},K_{3}^{\frac{2\sigma_{2}-1}{2-\sigma_{2}}}\right\}\max\left\{K_{2}^{\frac{3(1-\sigma_{1})}{2-\sigma_{1}}},K_{2}^{\frac{3(1-\sigma_{2})}{2-\sigma_{2}}}\right\},$
@@
$N(\sigma,T)$ $\leq\frac{1}{2\pi d}\left(\mathfrak{A}_{1}+\mathfrak{A}_{2}\cdot\left(\log{H_{\mathrm{RH}}}\right)^{-\frac{6(1-\sigma_{2})}{2-\sigma_{2}}}\right)T^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{T}\right)^{\frac{7-5\sigma}{2-\sigma}}$
@@
$N(\sigma,T)\leq 5.069\cdot T^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{T}\right)^{\frac{7-5\sigma}{2-\sigma}}+9.461\log^{2}{T}+167.8\log{T}$
@@
$N(T)\leq\frac{T}{2\pi}\log{\frac{T}{2\pi e}}+0.43\log{T}\leq 5.069\exp{\left(\frac{-0.56}{\frac{3}{2}-\frac{0.28}{\log{H_{\mathrm{RH}}}}}\right)}T\log^{2}{T}+9.461\log^{2}{T}+167.8\log{T}$ (11)
@@ 2.2. The fourth power moment estimate
Let $A_{0}\geq 10^{10}$. Then we have $\mathcal{M}_{2}(A_{0})\leq a_{1}(A_{0})$ and $\mathcal{M}_{2}(A_{0},T)\leq a_{2}(A_{0})T\log^{4}{T}$ for every $T\geq A_{0}$, where
$a_{1}(A_{0})\mathrel{\mathop{:}}=\frac{1}{2\pi^{2}}A_{0}\log^{4}\frac{A_{0}}{2}+34.842A_{0}\log^{3}\frac{A_{0}}{2}+1.3897\cdot 10^{16}$
@@
$a_{2}(A_{0})\mathrel{\mathop{:}}=\frac{1}{2\pi^{2}}\left(1-\frac{\log{2}}{\log{A_{0}}}\right)^{4}+\frac{34.842}{\log{A_{0}}}\left(1-\frac{\log{2}}{\log{A_{0}}}\right)^{3}+\frac{9.5079\cdot 10^{15}}{A_{0}\log^{4}{A_{0}}}.$
@@
The first estimate follows by the second inequality in (9) of Corollary 2. Because $\mathcal{M}_{2}\left(A_{0},T\right)\leq\mathcal{M}_{2}\left(10^{10},T\right)$ and $\mathcal{M}_{2}\left(10^{10},2\cdot 10^{10}\right)\leq 9.5079\cdot 10^{15}$ by Table 2, we have $\mathcal{M}_{2}(A_{0},T)\leq a_{2}(T)T\log^{4}{T}$ by (18), see the proof of Corollary 2. The derivative analysis shows that $a_{2}(T)$ is a decreasing function for $T\geq 10^{5}$, implying that $a_{2}(T)\leq a_{2}(A_{0})$. This gives the second estimate. ∎
@@
Assume $T\geq T_{0}^{\prime\prime}=T_{0}\geq T_{0}^{\prime}=3\cdot 10^{3}$, and also that the conditions from Theorem 2 are satisfied. Then
$\mathcal{M}_{2}\left(T_{0}^{\prime\prime},T\right)\leq\frac{1}{2\pi^{2}}T\log^{4}{\frac{T}{2}}+\mathfrak{F}_{1}\left(T_{0}^{\prime\prime},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)T\log^{3}{\frac{T}{2}}+\mathcal{M}_{2}\left(T_{0}^{\prime\prime},2T_{0}^{\prime\prime}\right)$ (18)
by (14). Also, $\mathcal{M}_{2}\left(T_{0}^{\prime\prime},2T_{0}^{\prime\prime}\right)\leq\mathcal{M}_{2}\left(T_{0}^{\prime},2T_{0}^{\prime\prime}\right)\leq\mathcal{C}_{2}\left(T_{0}^{\prime\prime}\right)$ and $\mathcal{M}_{2}\left(T_{0}^{\prime\prime}\right)\leq\mathcal{C}_{2}\left(T_{0}^{\prime\prime}/2\right)+7.032\cdot 10^{5}$ by (17) and (8), respectively, where
@@
Trivially, $\mathcal{M}_{2}(T)=\mathcal{M}_{2}\left(T_{0}^{\prime\prime}\right)+\mathcal{M}_{2}\left(T_{0}^{\prime\prime},T\right)$. We take the values for $T_{0}^{\prime\prime}$ as indicated in Table 3. Finding a global minimum for $\mathfrak{F}_{1}\left(T_{0}^{\prime\prime},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)$ for each $T_{0}^{\prime\prime}$ is a computationally very complex problem, since we need to optimize 11 parameters that satisfy the conditions (2)–(4) from Theorem 2. To simplify the process, we used randomization; we generate $50$ trials where for each trial we randomly choose $100$ sets of values for these parameters, then we select an element from this set such that $\mathfrak{F}_{1}\left(T_{0}^{\prime\prime},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)$ is minimal, and finally the “minimal” set from all trials. The values we obtained with such a method are listed in Table 3.
$T_{0}^{\prime\prime}$ $\bm{c}$ $\bm{\sigma}$ $\bm{\mathfrak{a}}$ $\bm{\mathfrak{b}}$
$10^{4}$ (0.879,0.1809)(-0.879,0.1809) $(0.3545,(0.7898,0.4008),(0.5456,0.3933))$ $(0.339,1.7553)$ $(1.7215,1.303)$
$10^{5}$ (0.938,0.194)(-0.938,0.194) $(0.368,(0.718,0.435),(0.629,0.3654))$ $(1.99,1.44)$ $(1.33,1.45)$
$10^{6}$ (0.9295,0.1721)(-0.9295,0.1721) $(0.4077,(0.6957,0.3413),(0.5672,0.4288))$ $(1.6769,1.5632)$ $(1.6891,1.6748)$
$10^{7}$ (0.9273,0.2061)(-0.9273,0.2061) $(0.4104,(0.6263,0.3536),(0.5743,0.41))$ $(1.8473,1.8288)$ $(1.649,1.9094)$
$10^{8}$ (0.9266,0.1637)(-0.9266,0.1637) $(0.3561,(0.7103,0.3381),(0.5645,0.4285))$ $(0.6164,1.9123)$ $(1.0018,1.9929)$
$10^{9}$ (0.834,0.22985)(-0.834,0.22985) $(0.3964,(0.9479,0.3927),(0.5846,0.4235))$ $(1.3609,1.1367)$ $(0.7112,1.8455)$
$10^{10}$ (0.8303,0.1676)(-0.8303,0.1676) $(0.3985,(0.668,0.4088),(0.551,0.386))$ $(1.4356,1.28)$ $(1.5563,1.6332)$
$10^{15}$ (0.813,0.221)(-0.813,0.221) $(0.3464,(0.8268,0.4263),(0.5698,0.3511))$ $(1.0851,1.0245)$ $(0.8137,1.8276)$
$10^{20}$ (0.875,0.1936)(-0.875,0.1936) $(0.3785,(0.8615,0.45),(0.619,0.393))$ $(0.1917,1.525)$ $(1.1358,1.717)$
Table 3. Our choice of the values for the parameters $\bm{c}$, $\bm{\sigma}$, $\bm{\mathfrak{a}}$ and $\bm{\mathfrak{b}}$.
We obtain $\mathfrak{F}_{1}\left(T_{0}^{\prime\prime},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)\leq\mathfrak{M}_{1}\left(T_{0}^{\prime\prime}\right)$, where the values for $\mathfrak{M}_{1}\left(T_{0}^{\prime\prime}\right)$ are listed in Table 2. The conditions (5) and (6) from Theorem 2 are clearly satisfied for each $T_{0}^{\prime\prime}$ by the above choice of parameters. The first inequality in (9) thus follows from (15) since $\mathcal{M}_{2}(T)\geq\mathcal{M}_{2}\left(T_{0}^{\prime\prime},T\right)$. Now, $\mathcal{M}_{2}\left(10^{4},2\cdot 10^{4}\right)\leq 9.9857\cdot 10^{8}$ and $\mathcal{M}_{2}\left(10^{4}\right)\leq 3.8095\cdot 10^{8}$, confirming the values listed in Table 2. This gives $\mathcal{M}_{2}\left(10^{4}\right)+\mathcal{M}_{2}\left(10^{4},2\cdot 10^{4}\right)\leq\mathfrak{M}_{2}\left(10^{4}\right)$, where the value for $\mathfrak{M}_{2}\left(10^{4}\right)$ is as in Table 2. The proof of (9) for $T_{0}=10^{4}$ is thus complete. For other values of $T_{0}$ we proceed as follows. For $n\in\{5,\ldots,10\}\cup\{15\}\cup\{20\}$, let $\mathscr{A}_{n}\mathrel{\mathop{:}}=\{1,\ldots,n-4\}$ if $n\in\{5,\ldots,10\}$, $\mathscr{A}_{n}\mathrel{\mathop{:}}=\{5,\ldots,11\}$ if $n=15$, and $\mathscr{A}_{n}\mathrel{\mathop{:}}=\{5\}\cup\{10,\ldots,16\}$ if $n=20$. Then
@@
$\mathcal{M}_{2}\left(10^{n},2\cdot 10^{n}\right)\leq\min\biggl\{\mathcal{C}_{2}\left(10^{n}\right),\\ \min_{i\in\mathscr{A}_{n}}\left\{\frac{10^{n}}{\pi^{2}}\log^{4}{10^{n}}+\left(2\cdot 10^{n}\right)\mathfrak{M}_{1}\left(10^{n-i}\right)\log^{3}{10^{n}}+\mathcal{M}_{2}\left(10^{n-i},2\cdot 10^{n-i}\right)\right\}\biggr\}$
and
$\mathcal{M}_{2}\left(10^{n}\right)\leq\min\left\{\mathcal{C}_{2}\left(\frac{10^{n}}{2}\right)+7.032\cdot 10^{5},\right.\\ \left.\min_{i\in\mathscr{A}_{n}}\left\{\frac{10^{n}}{2\pi^{2}}\log^{4}{\frac{10^{n}}{2}}+\mathfrak{M}_{1}\left(10^{n-i}\right)10^{n}\log^{3}{\frac{10^{n}}{2}}+\mathfrak{M}_{2}\left(10^{n-i}\right)\right\}\right\}.$
With this we recursively verify that $\mathcal{M}_{2}\left(10^{n}\right)+\mathcal{M}_{2}\left(10^{n},2\cdot 10^{n}\right)\leq\mathfrak{M}_{2}\left(10^{n}\right)$, where the values for $\mathfrak{M}_{2}\left(10^{n}\right)$ are listed in Table 2. The proof of Corollary 2 is thus complete. ∎
@@
$\mathcal{M}_{1}(T)\leq T\log{T},\quad\mathcal{M}_{2}(T)\leq\left(\frac{1}{2\pi^{2}}+\frac{\mathfrak{M}_{1}(T_{0})}{\log{T}}+\frac{\mathfrak{M}_{2}(T_{0})}{T\log^{4}{T}}\right)T\log^{4}{T}$
for $T\geq T_{0}\geq 10^{6}$ by [DZA22, Theorem 3.4] for the first inequality, and by Corollary 2 for the second inequality. The result now follows by the Cauchy-Bunyakovsky-Schwarz (CBS) inequality since then $\mathcal{M}_{2/3}(T)\leq\left(\mathcal{M}_{1}(T)\mathcal{M}_{2}(T)\right)^{1/2}$. ∎
@@ 3. Gabriel’s convexity theorem and applications
Let $1/2\leq\alpha<1<\beta\leq 3/2$, $T\geq H>1$, $\kappa>0$, $A\geq 0$, and also let $\lambda$ and $\mu$ be arbitrary positive numbers. Then
@@
Let $T\geq T_{0}\geq 2$ and $1/2\leq\alpha<\beta<1$. Assume that $f(z)$ is a holomorphic function in the strip $x=\Re\{z\}\in(\alpha,\beta)$ with continuous extension to the boundary, such that the conditions (2) and (3) from Theorem 4 are satisfied. Assume also that
@@ 4. Proof of Theorem 1
Let $T>0$, $X\geq X_{0}\geq 3\cdot 10^{10}$, and $0<\delta\leq\frac{1}{20}\log{X}$. Then
$F_{X}\left(1+\frac{\delta}{\log X},0,T,\frac{1}{2}\right)\leq\left(C_{3}(X_{0},\delta)+\frac{C_{4}(X_{0},\delta)\left(T+\pi m_{0}\right)}{X}\right)\log{X},$
@@
$C_{3}(X_{0},\delta)\mathrel{\mathop{:}}=\frac{\pi m_{0}b_{4}}{2\delta}\exp{\left(\frac{2\delta\gamma}{\log{X_{0}}}\right)},\quad C_{4}(X_{0},\delta)\mathrel{\mathop{:}}=\frac{b_{4}}{5\delta e^{\delta}}+\frac{b_{3}e^{-2\delta}}{\log{X_{0}}},$
with $b_{3}\mathrel{\mathop{:}}=0.605$, $b_{4}\mathrel{\mathop{:}}=0.54$, and $m_{0}$ as in Lemma 5.
@@
For $\tau\in(1,11/10]$ and $X\geq 3\cdot 10^{10}$ we have [^9] The second inequality through a private communication with Prof. Ramaré, since he and Zuniga Alterman had recently proved it and their result will be soon announced.
$\sum_{n=1}^{\infty}\frac{\lambda_{X}^{2}(n)}{n^{\tau}}\leq\sum_{n\leq X}\frac{1}{n^{\tau}}\left(\sum_{\begin{subarray}{c}d\mid n\\ d\leq X\end{subarray}}\mu(d)\right)^{2}+\sum_{n>X}\frac{\lambda_{X}^{2}(n)}{n^{\tau}}=\zeta(\tau)\sum_{d_{1}\leq X}\sum_{d_{2}\leq X}\frac{\mu(d_{1})\mu(d_{2})}{[d_{1},d_{2}]^{\tau}}\leq\frac{b_{4}e^{\gamma(\tau-1)}}{\tau-1},$
where $\lambda_{X}(n)$ are Dirichlet coefficients of $f_{X}(s)$. The proof now closely follows that of [KLN18, Lemmas 3.4 and 4.3], namely that
$F_{X}\left(\sigma,0,T,\frac{1}{2}\right)$ $\leq\pi m_{0}\sum_{n=1}^{\infty}\frac{\lambda_{X}^{2}(n)}{n^{2\sigma-1}}+\left(T+\pi m_{0}\right)\sum_{n=1}^{\infty}\frac{\lambda_{X}^{2}(n)}{n^{2\sigma}}$
$\leq\frac{\pi m_{0}b_{4}}{2}\frac{e^{2\gamma(\sigma-1)}}{\sigma-1}+\frac{T+\pi m_{0}}{X}\left(\frac{b_{3}}{X^{2(\sigma-1)}}+\frac{1}{5(5X)^{\sigma-1}}\sum_{n=1}^{\infty}\frac{\lambda_{X}^{2}(n)}{n^{\sigma}}\right)$
$\leq\left(\frac{\pi m_{0}b_{4}}{2}e^{2\gamma(\sigma-1)}+\left(\frac{(\sigma-1)b_{3}}{X^{2(\sigma-1)}}+\frac{b_{4}}{5(5X)^{\sigma-1}}e^{\gamma(\sigma-1)}\right)\frac{T+\pi m_{0}}{X}\right)\frac{1}{\sigma-1}$
for $\sigma\in(1,21/20]$ and $X\geq 3\cdot 10^{10}$. The proof now follows by taking $\sigma=1+\delta/\log{X}$ in the above inequality. ∎
@@
Let $\geq H\geq 2\pi m_{0}, TT$$\geq T_{0}>0$ and $X=hT$ for $h\geq 10^{9}/T_{0}$. Then
$\int_{H}^{T}\log{\left|h_{X}\left(\frac{3}{2}+\mathrm{i}t\right)\right|}\mathrm{d}t\leq C_{5}(hT_{0},h),$
@@
with $m_{0}$ as in Lemma 5.
@@
$\log{\left|h_{X}\left(\frac{3}{2}+\mathrm{i}t\right)\right|}\leq-\frac{\log{\left(1-b_{6}(hT_{0},1/2)^{2}\right)}}{b_{6}(hT_{0},1/2)^{2}}\left|f_{X}\left(\frac{3}{2}+\mathrm{i}t\right)\right|^{2}.$
@@
$-\frac{2\pi}{\log{2}}\log{\left(1-9\left(1+\frac{2+\gamma+\frac{7}{36u}}{\log{u}}\right)^{2}\frac{\log^{2}{u}}{u}\right)}.$ (34)
@@
Let $\geq H\geq H_{0}>1, TT$$\geq T_{0}\geq e^{e}$, and $\kappa>0$. Also, let $X=hT$ for $h\geq 3\cdot 10^{10}/T_{0}$, and let $w_{1}=w_{1}\left(\sigma,1/2,1+\delta/\log{X}\right)$ and $w_{2}=w_{2}\left(\sigma,1/2,1+\delta/\log{X}\right)$ for $0<\delta\leq\frac{1}{20}\log{(hT_{0})}$ and with $w_{1}$ and $w_{2}$ as in (21). Then
$F_{X}\left(\sigma,H,T,\frac{3}{4}w_{1}+\frac{1}{2}w_{2}\right)$ $\leq L_{0}^{\frac{2}{2-\sigma}}K_{3}^{\frac{2\sigma-1}{2-\sigma}}\left(L_{1}\cdot\left(K_{1}T\right)^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{T}\right)^{\frac{4(1-\sigma)}{2-\sigma}}\left(\log{(hT)}\right)^{\frac{1}{2-\sigma}}\right$.
$\left.+L_{2}\cdot\left(K_{2}T\right)^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{(hT)}\right)^{\frac{2\sigma-1}{2-\sigma}}\right)$
@@
$\int_{\frac{\kappa}{\mu}\left(\frac{A}{T}\right)^{2}}^{\infty}e^{-u}F_{X}\left(\beta,0,T\sqrt{\frac{\mu u}{\kappa}},\mu\right)\mathrm{d}u\leq K_{3}\log(hT)$